# Mixed Precision Iterative Refinement for Linear Inverse Problems

### v1

@@ 6 Conclusion
In this work, we have investigated IR on the Tikhonov problem in mixed precision. We showed that the iterates computed in mixed precision behave in a filtering manner by deriving methodology to formulate the iterates as a recursive relationship involving the iterates of preconditioned Landweber with a Tikhonov-type preconditioner and previous terms. Our numerical results showcase that simulated mixed precision IR can give comparable accuracy compared to results computed in double precision and superior results to the described AIR method which is appropriate and efficient to compute with in image deblurring applications.

### v2

@@ 6 Conclusion
In this work, we have investigated IR on the Tikhonov problem in mixed precision. We have shown that the iterates of IR on the Tikhonov problem computed in mixed precision behave in a filtering manner by deriving methodology to formulate the iterates as a recursive relationship involving the iterates of preconditioned Landweber with a Tikhonov-type preconditioner and previous terms. Our numerical results suggest that simulated mixed precision IR applied to inverse problems can give comparable accuracy compared to results computed in double precision and that solution variability when utilizing a low precision preconditioner can be influenced by the level of regularization as well as with the choice of the working and residual precisions. We also found that the MP-IR method to be superior when compared to the described AIR method which may be appropriate and efficient to compute with in image deblurring applications.
Acknowledgments
The authors wish to thank the anonymous referees for carefully reading the manuscript and providing many useful suggestions that improved the presentation of the work. Additionally, they would like to thank Alan Edelman and Evelyne Ringoot for providing their invaluable computing expertise. L. O. would like to thank Marco Donatelli for hosting a research visit to Como in 2025 which included fruitful discussions that also helped improve the manuscript. This work was partially supported by the National Science Foundation (NSF) under grants DMS-2038118 and DMS-2208294.